### v1

# Applications of Generative AI (GAI) for Mobile and Wireless Networking: A Survey

Abstract
The success of Artificial Intelligence (AI) in multiple disciplines and vertical domains in recent years has promoted the evolution of mobile networking and the future Internet toward an AI-integrated Internet-of-Things (IoT) era. Nevertheless, most AI techniques rely on data generated by physical devices (e.g., mobile devices and network nodes) or specific applications (e.g., fitness trackers and mobile gaming). To bypass this circumvent, Generative AI (GAI), a.k.a. AI-generated content (AIGC), has emerged as a powerful AI paradigm; thanks to its ability to efficiently learn complex data distributions and generate synthetic data to represent the original data in various forms. This impressive feature is projected to transform the management of mobile networking and diversify the current services and applications provided. On this basis, this work presents a concise tutorial on the role of GAIs in mobile and wireless networking. In particular, this survey first provides the fundamentals of GAI and representative GAI models, serving as an essential preliminary to the understanding of the applications of GAI in mobile and wireless networking. Then, this work provides a comprehensive review of state-of-the-art studies and GAI applications in network management, wireless security, semantic communication, and lessons learned from the open literature. Finally, this work summarizes the current research on GAI for mobile and wireless networking by outlining important challenges that need to be resolved to facilitate the development and applicability of GAI in this edge-cutting area.
Index Terms:
Artificial Intelligence, Generative AI, Internet of Things (IoT), Mobile Networking, Machine Learning, Wireless Networks.
I Introduction
Generative Artificial Intelligence (GAI), a prominent branch of Artificial Intelligence (AI) responsible for AI-Generated Content (AIGC), has recently attracted significant attention from the scientific community to big tech giants and governments. The three most popular representations of GAI are ChatGPT, Bing, and Google Bard, which are Large Language Model (LLM)-based chatbots, that can conversationally interact with humans and generate new data based on the input and contextual data [1]. Different from conventional AI and Machine Learning (ML) systems that are only trained on a certain dataset to create mapping abstractions to make highly accurate decisions, GAI platforms do not just learn from the input dataset but also rely on the interaction of data to combine, modify, and create new information to reinforce its learning process. Driven by consumer interests, GAIs have recently emerged as a powerful tool for various uses. For example, in robotics control, Reinforcement Learning (RL) is a key technique for enabling robots to learn from their own experiences, which are then translated into appropriate movements and actions in real time. Through GAI interaction with RL, robots become more humanoid in generating new motions and movement skills without human intervention or initial knowledge provision [2]. Another example is in the field of cybersecurity for the Internet-of-Things (IoT), Deep Generative Models (DGM), fundamental components behind GAI, help generate patterns and address critical challenges involving the lack of malicious samples and the rapid evolution of cyberattacks [3]. As a consequence, the adoption of GAIs is projected to produce new ways to significantly improve current AI-enabled algorithms in mobile networking and the IoT.
Figure 1: Illustration of a high-level GAI architecture.
Fig. 1 illustrates a high-level GAI architecture, where the paradigm is composed of three main blocks. The “grey” block refers to the GAI models, in which the selection and optimization depend on use cases, relevant data, and performance indicators. Pretrained LLMs and Generative Adversarial Networks (GANs) are two representative GAI models that have found a lot of success in real GAI-enabled applications, such as chatbots, code generation, search engines, and cybersecurity. The “orange” block indicates the system settings used as the input to the GAI models to generate the output data. More precisely, users can provide initial guidance about the data to be generated and the links between the input data and expected output. For example, a typical user can input “what is generative AI?” and “what are essential blocks of GAI?” if they seek to receive draft answers from GAI tools. Finally, the “green” block corresponds to data generated from GAI models based on input from users (e.g., text, audio, video, and code), which then are used to provide other services, such as enhanced network management and improved wireless security in mobile communication and wireless networks. Fig. 2 shows a comprehensive overview of popular GAI developer tools and services, showcasing the layered structure from foundational hardware to essential cloud services and advanced model safety tools.
Figure 2: Popular developer tools and services for building GAI frameworks.
I-A State-of-the-Arts
AI techniques have become widespread in various research areas in mobile networking and the IoT, such as network management, wireless security, resource allocation, and edge services. In recent years, data-driven and AI-enabled approaches have established clear performance superiority over many conventional optimization-based designs, which are limited primarily by network scalabilities in handling massive connections and heterogeneous variations [4]. Additionally, AI techniques enable efficient approximation of computationally challenging algorithms, to effectively overcome nonlinear and unknown patterns in wireless networks, and rapidly accelerate existing model-based algorithms. On this basis, there have been several surveys of different AI paradigms and their applications for mobile networking, such as Deep Learning (DL) [5, 6], Deep Reinforcement Learning (DRL) [7, 8], Federated Learning (FL) [9, 10], and swarm intelligence [11]. However, current AI-based approaches for mobile and wireless networking face several issues. Specifically, conventional AI algorithms usually require a large volume of labeled training data to facilitate better performance. Unfortunately, collecting and labeling training data in complex environments like wireless networks are costly, time-consuming, and require intensive human interventions. In addition, with an enormous number of diverse users, future wireless systems are highly dynamic and uncertain. It can significantly degrade the performance of conventional AI models, which are usually trained in a specific environment/condition. Although the distributed learning paradigm can help to facilitate the learning process of AI-based approaches, it still suffers heavily from scarcity of data required for efficient training [12]. Moreover, due to the differences in local system characteristics, the distributed AI system suffers from the long-tailed distributed data phenomenon. This data imbalance results in a negative transfer toward the data cluster with a dominant size. To address all these practical challenges, GAI is considered an efficient solution thanks to its properties of domain adaptation, data generation, and abnormal detection. Aiming to address these challenges, there are several surveys in the literature focusing on the applications of GAI in various domains such as mobile networking, Semantic Communication (SemCom), and security and privacy, as summarized in Table I.
For example, GANs, a notable GAI technique, are reviewed in several recent works [13, 14, 15, 16, 17, 18], providing preliminary discussions about the use of this GAI-specific technique in emerging networks. More recently, GAI has been reviewed for several research issues in wireless networks. For instance, in [19], the authors provide a detailed survey of training solutions for GANs that deal with challenges of model collapsing and gradient vanishing, providing meaningful techniques and guidance for mobile network deployments. Similarly, in [20], the authors analyze GANs with various stability issues and survey several training solutions by assessing the obstacles. In [21], several critical GAN analyses of computer and communication networks are outlined, with a focus on network management, security, and mobile networks. Differently, the authors in [22] introduce the concept of on-device GAI, where multiple devices, functioning as distinct learning agents, deploy on-device GAI models for the collaborative processing of complex tasks. In [23], a direction on GAI services at the network edge has been outlined in various use cases, followed by guidance on building a hierarchical collaboration architecture of IoT devices, mobile edges, and centralized clouds. In [24], wireless network management efficiency can be improved by DL models in general and particular applications of DGM. In addition, the authors in [25] present a review of the applications of GAI for physical layer communications, focusing on emerging problems such as signal detection, physical layer security, and Joint Source-Channel Coding (JSCC). Considering a different perspective, the authors in [26] discuss the potential of common GAI techniques, e.g., GANs, VAEs, and normalizing flows, in addressing the design challenges of unmanned vehicle swarms.
On another front, the interplay of GAI and SemCom has opened up a new avenue for reducing resource consumption while maintaining the reliability of wireless communication networks [27]. Fundamentally, SemCom relies on two fundamental techniques [28, 29] of knowledge extraction and lossy data compression, both rooted in AI. As such, the current landscape of SemCom designs represents a paradigm shift, with these innovative concepts emerging across various disciplines. In the literature, several surveys aim to elucidate the intricate connections between the potential AI techniques within GAI and their application in the field of SemCom. From the perspective of Knowledge Graph (KG), the current SemCom is based on the KG embeddings from Natural Language Processing (NLP), which are comprehensively reviewed in [30, 31, 32, 33]. From the perspective of lossy compression, despite a huge range of research that has both integrated and without knowledge representation to improve data reconstruction performance, no studies have surveyed these techniques in general and lossy compression-aided SemCom in particular.
In addition to the aforementioned dominant features of GAI, it also possesses significant merits in the cutting-edge area of security. For instance, in [34], Zhang et al. explore GANs in transportation, including autonomous driving, data completion, and traffic anomaly detection. In [35], Liu et al. explore how adversarial ML can be both a threat and a defense mechanism in mobile networks. In [36], Cai et al. provides a discussion of the potential benefits and challenges of GANs in the context of privacy and security in mobile networks. In [37], Deldjoo et al. present the role and application of GANs in improving data distribution learning of recommendation systems. In [38], Dunmore et al. survey and discuss the potential of GAN applications in enhancing cybersecurity in mobile networks, particularly in Intrusion Detection Systems (IDS). In [39], Lin et al. investigate and inspect the use of GANs in enhancing trajectory prediction, security, and anomaly detection of the wireless-based transportation domain. However, a recent investigation in [40] points out that the use of GAI exposes potential security vulnerabilities, and this will generate a double-edged sword for wireless security if GAI is not exploited properly.
I-B Motivations and Contributions
A common idea in existing studies is that most focus is on clarifying either applications or challenges. The increasing emphasis on generative AI technologies is expanding into various applications and domains within mobile and wireless networking. However, current studies do not provide comprehensive GAI evaluation and its applications. To the best of the authors’ knowledge, the fundamentals and applications of GAI as outlined above have not been thoroughly examined in the literature. Consequently, we aim to bridge this gap by providing the necessary preliminaries to GAI and its key applications in the future of the IoT and mobile networking. This work centralizes all recent developments in GAI and complementary technologies and remarkable advancements in various aspects of mobile networks. At the same time, this work also seeks to gather the recent works in this trending research field and provide a unified repository. The notable contributions of our work are listed below.
  • We present an overview of GAI, covering its historical development, foundational structures (e.g., autoencoders, GANs, and diffusion models), data creation, and multimodal aspects (e.g., meta-learning and multi-tasking).
  • We summarize the contributions of GAI to network management, particularly in anomaly detection and security through its ability to learn and flag deviations.
  • We explore the applications of GAI in wireless security based on baseline models, detecting anomalies, fortifying against cyber threats, improving intrusion detection, and optimizing security policies.
  • We examine the applications of GAI in enhancing SemCom through NLP for human-like text, chatbots, virtual assistants, automated content creation, and improved language translation.
  • We reveal challenges of GAI in mobile networking, emphasizing technical complexities, legacy systems, ethical concerns, security, and privacy and necessitating comprehensive solutions.
Table I: Existing survey papers about GAI for mobile networking. ✓, ✗, and ✱  indicate that the topic is well-covered, uncovered, and partially covered, respectively.
Ref. Year Focus ML GAI NetMan Security SemCom
[13] 2019 Autonomous wireless channel modeling without complex theoretical analysis.
[14] 2022 Assistance of GAN in anomaly detection, with analysis techniques and application domains.
[15] 2022 Generative adversarial learning for secure, and reliable trust management in real-time communications.
[16] 2023 AIGC for efficient and accessible content creation using large-scale models with unimodal and multimodal interactions.
[17] 2023 Focus on AIGC model architecture to output types, in tasks like text, images, and videos.
[18] 2023 Impact of GAI on different industries and taxonomy of recent key models.
[19] 2021 Surveying training solutions for GAN with mode collapse, and gradient vanishing.
[20] 2021 Analyzes GANs with various stability issues, and training solutions by assessing obstacles.
[21] 2021 Explores aspects of computer and communication networks, including network management, security, and mobile networks.
[22] 2023 Integration of multi-agent GAI in wireless networks based on cloud-based LLMs.
[23] 2024 Personalized AIGC services at mobile edge networks ensuring privacy measures.
[24] 2023 Applications of deep generation models in improving the efficiency of wireless network management.
[27] 2023 Architecture, challenges, and outlook of Generative AIs.
[34] 2019 Explores GAN in transportation including autonomous driving, data completion, and traffic anomaly detection.
[35] 2021 Adversarial ML in autonomous driving, data generation, and anomaly detection.
[36] 2021 Focuses on GAN for privacy and security contexts, highlighting potential benefits and challenges.
[37] 2021 Security of recommender systems and its application in GANs for data distribution learning.
[38] 2023 GANs in IDS and their applications in cybersecurity research.
[39] 2023 Trajectory prediction, security, and anomaly detection in the transportation domain.
[41] 2023 Generative diffusion-guided framework for SemCom, using diffusion models.
[42] 2023 Causal SemCom for digital twin-based wireless systems.
[43] 2023 FL solutions to enhance the efficiency and privacy protection of AIGC using stable diffusion models.
[44] 2023 AaaS architecture that deploys AIGC models in edge networks.
This Work - A review on applications of GAI for network management, security, and SemCom in mobile networking.
I-C Paper Organization
In the study, Section II presents a preliminary to GAI, including the development history, data-processing and data-creation techniques, and representative models. In Section III, we comprehensively review the role of GAI in mobile network management. Section IV reviews the use of GAI in various wireless security issues, including data obscurity, intrusion detection, jamming attacks, and generative steganography, while Section V describes the applications of GAI for research advances in SemCom. Section VI discusses crucial challenges and explores potential research directions to facilitate applications of GAI techniques. We finally conclude the paper in Section VII.
II Generative AI: Preliminary
This section introduces a preliminary to GAI, along with its structure and data-creation techniques. It lays the platform for building robust applications of GAI in mobile networking and the future Internet.
II-A History of Generative AI
AI has a long history and has developed a lot over many years. Due to big investments from governments and large companies, AI has reached important milestones and is used in almost every industry, including infotainment, finance, advertising, defense, agriculture, and telecommunications.
In general, AI techniques and their applications are designed by researchers in the field of mobile networking. AI-driven algorithms are used by researchers to analyze and optimize data traffic, improving network efficiency and performance, while automating network management tasks like configuration and troubleshooting [45]. These algorithms also facilitate the dynamic allocation of network resources based on real-time demand and usage patterns, optimizing resource utilization and enhancing network capacity and scalability [46]. To ensure network security, AI-powered IDS analyzes network traffic patterns in real time, identifying and mitigating potential threat, and supporting network resilience against cyber attacks and unauthorized access [47]. Additionally, AI-based encryption and authentication mechanisms ensure data confidentiality and integrity, protecting sensitive information transmitted over mobile networks from unauthorized access and tampering [48].
Table II: Advantages and challenges of GAI use cases based on its characteristics.
Ref. Characteristics Description Advantages Challenges
[49] Noisy data Handles uncertainty in data generation to maintain the stability of results. • Generated diverse and variable data. • Recovered data during training. • Inconsistent output quality. • Unexpected results.
[50] Data augmentation Improves training datasets by generating additional data samples to better train the model. • Increased size and diversity of the training dataset. • Overfitting reduction. • High computational resources and time for generating augmented data. • Instability in model improvement.
[51] Transfer learning Leverages pre-trained models to generate new data for specific tasks without full retraining. • Improved time and resources by reusing learned representations. • Useful in fine-tuning models for new tasks efficiently. • Limited to tasks similar to the pre-trained model and entirely different domains. • Faces with domain shift or bias issues.
[52] Style transfer Applies the style of one dataset to another, e.g., changing the artistic style of images. • Facilitates artistic expression by transforming content while preserving style. • Useful in image and video editing and creative applications. • Artifacts or distortions appear during style transfer. • Preserving content details can be challenging for complex styles.
[53] Text-to-image Generates images from textual descriptions, enabling AI-powered content creation. • Opens possibilities for generating images based on textual ideas or concepts. • Aids in creating visual content from text for various applications. • Difficulty in generating highly detailed images from textual descriptions. • Limited by the textual context and language-understanding capabilities.
[54] Super resolution Enhances image resolution, useful in improving image quality for various applications. • Enhances image quality for better visual perception and recognition. • Useful in medical imaging, surveillance, and content restoration. • Increased computational complexity and resource requirements. • May not always achieve satisfactory results for all types of images.
[55] Music generation Creates music compositions, often employing recurrent neural networks for sequence modeling. • Facilitates music composition and creativity in the absence of human composers. • Provides a wide range of music styles and genres. • Difficulty in generating music with the same level of emotional expression as human composers. • May produce music that lacks coherence or structure.
[56] Video synthesis Generates videos by predicting frames sequentially, enabling deepfake applications. • Enables video synthesis for various utilities, including entertainment and training. • Allows video generation based on textual descriptions. • Raises ethical concerns regarding fake videos or misinformation. • High computational and data requirements.
[57] Anomaly detection Identifies anomalies in data, useful for detecting outliers or unusual patterns. • Enhances the detection of rare and critical events in large datasets. • Useful for security and fault detection in various industries. • Challenges in labeling anomalies for training, leading to imbalanced datasets. • Risk of false positives and negatives based on the anomaly definition.
[58] Data privacy This raises concerns about potential misuse, as GAI can create realistic fake content. • Raises awareness about data privacy and the need for robust detection of fake content. • Encourages the development of countermeasures against misuse. • Potential for generating deepfake content that can be used maliciously. • Challenges in distinguishing between genuine and fake content in some cases.
Despite showing great potential in various domains, conventional AI approaches may not be effective in highly dynamic and complex environments (e.g., future mobile and wireless networking), as discussed in Section I. To overcome this drawback, the potential of GAI has recently been demonstrated by advances in VAEs [59] and GANs [60]. These innovations were based on Deep Neural Network (DNN) architectures capable of learning generative models for complex data types like images. In 2017, Google introduced the Transformer architecture [61] that further enhanced the capability of GAI in learning generative models. This led to the development of the first Generative Pre-trained Transformer (GPT), as known as GPT-1, in 2018. Since then, GAI has received attention from both academia and industry, with various advantaged techniques and applications, e.g., DALL-E and GPT-4. The advantages and challenges of GAI use cases can be summarized in Table II. In the following, we present the fundamentals of common GAI techniques together with their potential in mobile and wireless networking areas.
II-B Primitive Structure of GAI
One of the core objectives of generative adversarial learning-enabled trust management services for Beyond Fifth-Generation (B5G)/Sixth-Generation (6G) wireless networks is to enhance intelligence and security while ensuring reliable and real-time network security and service quality optimization [15]. The core elements inherent in the basic structure of GAI models empower them to provide robust and reliable services for both modern wireless service providers and users.
II-B1 Variational Autoencoders
Variational Autoencoders (VAEs), illustrated as in Fig. 3, play a crucial role in GAI by facilitating the generation of realistic and high-quality data. They are applied across diverse domains like Computer Vision (CV), NLP, and mobile networking. Especially in situations with limited labeled data, autoencoders are highly valuable as they can generate synthetic data samples for training purposes [62]. Furthermore, VAEs contribute significantly to dimensionality reduction and feature extraction, leading to a more compact and structured representation of complex data. Fundamentally, the encoder compresses data into a lower-dimensional representation, namely latent space, while the decoder module reconstructs the data from the latent space, restoring it to the original form. Once trained, the decoder becomes capable of generating new data points by sampling from the latent space [63]. This remarkable capability makes VAEs highly valuable for tasks such as data synthesis, text generation, anomaly detection, and wireless networks.
Figure 3: An illustration of an autoencoder-based GAI. Here, the generated data is guaranteed to be robust to noise as the hidden layers are being injected by noise during the training to simulate the channel noise of the wireless channel.
II-B2 Diffusion Models
An alternative approach to generative modeling is offered by diffusion models, which draw inspiration from non-equilibrium thermodynamics [64]. These models use a Markov chain of diffusion steps to gradually add random noise to the input data. Then, they learn to reverse this diffusion process to create desired data samples from the noise. Unlike VAE or flow models, diffusion models use a fixed training procedure and have high-dimensional latent variables that match the dimensionality of the original data. Table III highlights the advantages and limitations of GAI models in mobile networking applications.
GAI models are intended to generate new data samples that resemble a given training dataset. A GAN trained on a dataset of landscape images, for example, can produce new landscape images that resemble the training data but are not exact replicas of any particular image [65]. In GAI, there are several methods to produce data, such as likelihood-based models, energy-based models, and two-star models [17]. Estimating the probability distribution of the training data is the main goal of likelihood-based models, such as the Gaussian Mixture Model (GMM). GMM assumes that the data are produced by combining different Gaussian distributions. Data collection entails selecting samples from the learned distribution via a trained GMM. For instance, we can train a GMM with a handwritten number dataset; after learning its features, the data can be configured to produce novel samples [66].
Energy-based models define an energy function that assigns low energy to low-resolution samples from the training set and high energy to other samples. To find low-energy samples that are probably similar to the training data, one can adopt Markov Chain Monte Carlo (MCMC) techniques to search the sample space [67].
An alternative method involves incorporating GANs into diffusion models. In this approach, the generator endeavors to mislead the discriminator, while the discriminator strives to accurately classify the samples. To achieve this, these two models are trained in opposition to each other [68].
II-B3 Generative Adversarial Networks
GANs [60] were introduced to generate realistic synthetic data. The key rationale behind GANs lies in their unique architecture, which consists of two neural networks: the generator and the discriminator, engaged in a competitive game. The discriminator distinguishes genuine and generated data, while the generator creates synthetic data samples that closely resemble real data. This interplay drives an adversarial training process where the generator aims to deceive the discriminator and enhance the quality of its output. GANs have diverse applications in generative AI, including text generation, data augmentation, and image synthesis. They excel in generating a wide range of samples and capturing complex data distributions. Furthermore, the incorporation of techniques such as conditional GANs [69, 70], style transfer [71, 72], and progressive training [73] enhance the control and manipulation of the generated outputs. These advancements have made GANs indispensable in various industries, including art, fashion, and entertainment. Such methods also facilitate the production of diverse and realistic data, where their adversarial training process can capture complex data distributions and adapt to different applications. The use of GANs has transformed the field of GAI, enabling the creation of unique and high-quality content across multiple domains.
Table III: Advantages and Limitations of GAI Models in Mobile Networking Applications.
GAI Model Advantages Limitations
Autoencoder • Efficient data compression for mobile devices with limited storage. • Low-latency encoding and decoding suitable for real-time mobile applications. • Privacy preservation through on-device data processing. • Unsupervised learning without labeled data. • Well-suited for image reconstruction tasks. • Limited capacity to handle diverse and dynamic mobile data. • Sensitivity to variations in input data quality, affecting reconstruction. • High computational load during training on resource-constrained mobile devices. • Limited diversity in generated samples. • Sensitive to input noise and may suffer from overfitting.
Diffusion Models • Effective in modeling sequential dependencies in mobile data. • Privacy-preserving data generation through controlled diffusion. • Robust handling of diverse mobile data types. • Strong theoretical foundation with tractable likelihood. • Stable training without mode collapse. • Computationally demanding, challenging for mobile deployment. • Potential issues in capturing long-range dependencies in sequential data. • Training complexity increases with the scale of mobile datasets. • Sequential generation can be computationally expensive. • May struggle with capturing long-range dependencies and handling discrete data.
Generative Adversarial Networks • Creation of realistic synthetic data for diverse mobile applications. • Enhanced mobile security through adversarial training against attacks. • Suitable for mobile gaming and augmented reality applications. • High-quality sample generation in various domains. • Effective in learning complex data distributions. • Training instability might lead to mode collapse. • Prone to generating biased data, impacting fairness in mobile applications. • Resource-intensive training and generation process on mobile devices. • Lack of explicit likelihood estimation. • Sensitivity to hyperparameter tuning.
II-C Multi-modal GAI
Multi-modal GAI encompasses a range of tasks, including but not limited to image recognition, semantic segmentation, NLP, speech synthesis, and emotion detection. Through the integration of diverse modalities like visual, auditory, and textual information, multi-modal GAI is capable of capturing a holistic representation of the world, mirroring the way humans perceive and interpret multiple sources of information concurrently [74]. In what follows, we explain how GAI deals with multi-tasking problems.
II-C1 Meta-learning GAI
Using meta-learning [75], GAI systems can achieve fast adaptation by leveraging the task-specific gradient trajectories. This approach enables GAI systems to acquire knowledge more efficiently and transfer it across diverse domains [76], while also possessing the capability to continue learning [77, 78]. Additionally, meta-learning addresses the challenge of continuous learning by mitigating the issue of catastrophic forgetting, allowing GAI to retain previously acquired knowledge while continuously acquiring new knowledge. For example, the authors in [79, 80] propose a data acquisition framework based on GANs and meta-learning to reduce the need for collecting a large number of real samples. This framework generates wireless data samples for DL models with the same performance as models trained with more real samples. The research in [81] proposes a different way of leveraging meta-learning into GAI design. Instead of applying meta-learning to quickly adapt the GAI model from different domains, MetaMask leverages meta-learning to efficiently adapt the joint GAI model via trajectory estimation from the GAI sub-model. Fig. 4 shows the taxonomy of meta-learning and its application using GAI.
Figure 4: Taxonomy of Meta-learning through GAI.
II-C2 Multi-task GAI
Multi-task GAI refers to the ability of an AI system to effectively perform and generalize across a wide range of tasks [82, 83, 84]. By leveraging the knowledge and skills gained from one task and applying them to related or similar tasks, multi-task GAI utilizes transfer learning to improve performance across various domains [85, 86]. Moreover, these systems extract shared representations and beneficial characteristics that contribute to improved performance on new and unexplored tasks. Collaborative learning from multiple tasks reduces training time and enhances computational efficiency, optimizing resource utilization. Another advantage of multi-task GAI is domain adaptation, enabling the system to adjust its knowledge and skills to different domains or tasks [87]. This adaptability empowers the system to achieve high performance in unfamiliar or dynamic environments. Additionally, multi-task GAI facilitates continuous learning by gradually introducing new tasks while minimizing the forgetting of previously mastered tasks. This dynamic approach allows ongoing knowledge upgrades and long-term proficiency across diverse disciplines.
II-D Representative GAI Models
Across a diverse range of tasks and domains, several representative GAI models have demonstrated impressive capabilities. One such example is GPT-3, a renowned model developed by OpenAI. GPT-3 stands out for its exceptional language processing abilities, excelling in tasks like language generation, translation, summarization, and question-answering [88]. In addition, OpenAI’s DALL-E model is a specialized form of GAI that focuses on generating images from textual descriptions, effectively transforming natural language prompts into accurate visual representations [89]. It enables tasks such as image categorization, object detection, and text-based image synthesis. Contrastive Language-Image Pre-training, another creation from OpenAI, integrates language and image understanding in a cohesive manner [90]. These GAI models exemplify the potential for developing intelligent systems that excel across various activities and domains, showcasing cutting-edge progress in the field of AI.
III Application of Generative AI in Network Management
As modern networks exponentially grow in complexity and scale, traditional management approaches face challenges in ensuring efficiency, security, and adaptability. Integrating GAI intro mobile networks help overcome these circumvents, approaching the era of intelligent network management. To explain how GAI works, this discusses its particular applications from network management, network routing, and resource allocation to channel estimation. Then, we conclude this section with useful learned and key takeaways.
III-A Network Management
Software-Defined Networking (SDN) represents a pivotal convergence point between GAI and network management. To be specific, SDN’s separation of the control plane and data plane facilitates centralized network management and programmability. Integrating GAI into SDN frameworks empowers network administrators with enhanced automation, intelligent resource allocation, and adaptable network optimization. Specifically, GAI is in charge of real-time decision-making for effective traffic routing, load balancing, and Quality-of-Service (QoS) management, enabling SDN controllers to dynamically learn from and adapt to changing network conditions. For example, DGMs are used in [24] to address the issues of data scarcity, limited flexibility, and dynamic network states using a diffusion model to generate effective contracts. In [91], NetGPT is proposed to employ LLMs at the edge and the cloud in conjunction with location-based information to speed up personalized procedures and cloud interaction, enhancing intelligent network management and orchestration.
Network slicing, a vital aspect of network management, has great potential to leverage GAI and opens a new chance to customize network instances within a shared infrastructure [92]. On this basis, GAI can exploit this property to perform network data analysis, pattern prediction, and generation of optimized network slices tailored to specific application requirements. In this way, mobile networks can dynamically allocate resources, balance workloads, and adapt to changing traffic demands, contributing to better user experiences, increased network efficiency, and greater flexibility in service provisioning [93].
III-B Resource Allocation
In the context of dynamic resource allocation, GAI contributes to dynamically changing network conditions in real time through continuous monitoring and analysis of network data. This enables efficient utilization of resources, ensuring optimal performance and responsiveness to the evolving needs of various network factors such as network load, user demands, and QoS demands. It also helps systems use proper resource allocation strategies to ensure an optimal allocation that aligns with network objectives and user expectations.
For instance, the authors in [94] propose GAI-based resource allocation for the SemCom framework based on YOLOv7-X. This method helps reduce costs and improve information transfer in communication services by extracting the necessary semantic information from edge device images before deciding on resource allocation via confidence and AI-generated models. In [44], an AIGC-as-a-Service (AaaS) architecture is introduced to wireless edge networks, where GAI at edge servers is responsible for repairing damaged images and naturally creating high-resolution/augmented reality/virtual video content, reducing computational resources. Compared to DRLs, this architecture has not only a simple transmission protocol, but also an efficient and scalable method for mobile users to access intelligent services on demand without limited resource concerns.
III-C Network Routing
To deal with the dynamics and uncertainty of network conditions, introducing GAI into network routing can be an innovative alternative to conventional AI approaches. Specifically, GAI models can assess the efficacy of various routing algorithms and optimize routing choices for certain network objectives by simulating and creating synthetic network scenarios. This guarantees effective resource utilization, low latency, and better overall network performance. Moreover, by analyzing network data to spot potential weaknesses and learn from historical security incidents, GAI helps navigate and enhance routing patterns to reduce the possibility of assaults or unauthorized access, proactively mitigating security concerns.
A typical example of this feature is the adoption of DGMs in network economics to improve the efficiency of wireless network management by exemplifying and generating effective contracts for incentivizing mobile AIGC services [95]. Another example of GAI in routing solutions for different network status distributions and topology is exploiting a transfer RL algorithm to improve training efficiency by rapidly transferring knowledge [96].
III-D Channel Estimation
Numerous benefits can be achieved by incorporating GAI approaches into mobile networking channel estimation. First, GAI enhances the ability to detect faulty signals and provide appropriate fixes to improve channel estimation procedures. Second, GAI can streamline the resources used for channel estimation, which lowers the computational complexity and energy usage of conventional estimation techniques. Third, by learning the fundamental patterns and properties of wireless channels, GAI makes channel estimation more reliable and precise, particularly in difficult situations with little training data or time-varying channel circumstances.
To verify these promising traits, the work in [97] designs a high dimensional channel estimation approach based on a deep generative network combined with a compressive sensing strategy. The experiments from [97] indicate that GAI not only uses fewer channel pilots than conventional techniques, but also improves the execution time of channel estimation procedures for one-bit quantized pilot measurements. Similarly, another study for wideband channel estimation [98] has demonstrated the efficacy of GANs in dealing with channel estimation in the conditions of limited channel pilots and low transmit SNR. Specifically, it significantly reduces the number of pilots by about 70%, even at mmWave and THz frequencies, without affecting the estimation errors. On another front, the work in [99] displays the potential of deep score-based generative models to multi-input multi-output systems, with high fidelity channel estimation for channel communication of sizes 64×256 with pilot density up to 25%.
III-E Lessons Learned and Keys Takeaway
Modern networks have undergone a revolution because of the use of GAI in network management, which has made it possible to provide accurate and high-quality data. Among GAI techniques, GANs have become popular in many applications, such as picture synthesis, text generation, and data augmentation, providing innovative and varied training data for AI models. In wireless networks, GAI is crucial for allocating resources effectively, optimizing input data to cut transmission costs and increase accuracy. Additionally, generative models have solved practical issues in channel estimation by identifying network risks with greater precision and eliminating data imbalance problems.
The landscape of network routing presents challenges, encompassing aspects like security, user verification, and data integrity. Recent research indicates that the GAI is instrumental in addressing network routing challenges. It excels in enhancing security, conducting user verification, ensuring secure communication, and fortifying against advanced attacks. Future investigations should not be confined to refining GAI algorithms for optimized mobile network operations. Instead, they should focus on fortifying data security and continually enhancing user experiences. Moreover, the dynamic potential of GAI in mobile networking requires ongoing exploration for improved creativity and security in communication networks.
In summary, the key takeaways from this section are as follows.
  • GAI enhances network management by employing predictive models to optimize traffic flow and minimize downtime.
  • By dynamically adjusting computing resources based on real-time demand and ensuring optimal efficiency, the resource allocation capabilities are improved.
  • Channel estimation and security benefit are achieved from generative models with improved estimation and detection accuracy.
  • GAI can strengthen network security utilizing advanced anomaly detection algorithms to identify and mitigate potential threats.
  • GAI can contribute toward data confidentiality and integrity by implementing encryption protocols and monitoring data for unauthorized access or alterations.
IV Application of GAI in Wireless Security
Wireless security is a critical concern in mobile networking, with the ever-growing need to protect sensitive data and ensure secure communication. In recent years, GAI has emerged as a powerful tool with diverse applications, including anomaly detection, intrusion detection, authentication, and encryption. By harnessing the capabilities of generative models, novel approaches are being developed to enhance the resilience and effectiveness of security mechanisms in mobile networks. This section delves into the advancements, challenges, and potential directions in leveraging GAI for wireless security, contributing to the evolving landscape of mobile networking.
IV-A Sensitive Data Obscurity
Data sharing has become a popular trend in intelligent systems for training machine learning models. However, this activity poses numerous risks related to user privacy, in that sensitive information may be disclosed by third parties without unauthorized access or interception [100]. In this context, GAI has recently demonstrated its promising capabilities in obfuscating sensitive data by creating synthetic data that closely mimics real datasets [101].
There are multiple methods for generating synthetic data using GAI models, such as using GAN, VAE, or GPT models. In particular, data perturbation is one of GAI’s main techniques in safely hiding sensitive data [102]. This technique uses synthetic data to complicate extraction of useful information, by applying controlled perturbations to sensitive data [103]. In addition, distributing and transforming only the sensitive elements in datasets can also increase the security of data sharing while preserving the features and statistical utility of the original data for model training activities [104]. Moreover, GAI models can apply other techniques like differential privacy and data anonymization.
GAI techniques also aid in encrypted communication. Indistinguishable noise can be created from encrypted data using generative models, complicating decryption of the original data without the proper decryption keys [105]. With this strategy, sensitive data are sent more securely over wireless networks while maintaining their confidentiality and integrity. However, despite the apparent advantages, there are issues when using GAI to obfuscate sensitive data, such as the trade-off between data value and obscurity, the scalability of generative models of huge datasets, as well as the criteria and requirements for the reliability and efficacy evaluation of data obscurity strategies.
IV-B Intrusion Detection
Intrusion detection in wireless security is essential for identifying and combating threats or intrusions on mobile networks. Applying GAI techniques opens new avenues to strengthen IDSs and enhance wireless network security as a whole [106]. A valuable tool for discovering anomalies and distinguishing malicious activity in wireless network traffic is provided by GAI models like GANs. These models are capable of learning the typical patterns and behaviors of network traffic through training on large-scale datasets, allowing them to detect unforeseen attack patterns from expected norms with high accuracy [107].
Generating synthetic data samples that closely resemble actual network traffic is one of the main benefits of utilizing GAI in intrusion detection. IDS is capable of accomplishing this by enhancing its training data and producing a variety of testing and evaluation situations. IDS can increase its robustness and adaptability to various sorts of threats by creating synthetic attack patterns to improve its detection capabilities. The early discovery of zero-day attacks, vulnerabilities or exploits of which security experts are unaware, can also be helped by GAI [108]. IDS can detect suspicious actions that depart from usual behavior by continually monitoring network traffic and analyzing patterns using generative models, acting as an early warning system for potential zero-day attacks.
Furthermore, the effectiveness and precision of IDS can be increased using GAI approaches. IDS can improve their detection accuracy and minimize the impact of false alarms by lowering false positives and false negatives [109]. Additionally, the use of generative models can lower the computing burden of IDS, enabling instantaneous network traffic monitoring and analysis [110]. However, integrating GAI into IDSs can be difficult. The necessity for extensive wide-ranging training datasets, the interpretability of generative models, and the adversarial robustness of the models against complex attacks are a few complications.
IV-C Jamming Attacks
The integrity and availability of mobile networks are seriously threatened by jamming assaults in the area of wireless security. These assaults encompass deliberate disruption or interference of wireless signals, which results in a reduction in service or prevention of communication [111]. Innovative methods for countering and lessening the effects of jamming attacks in mobile networking are possible using GAI techniques. GANs can be useful in recognizing and preventing jamming attacks. These models can recognize irregularities brought on by jamming activities because they can learn and comprehend the typical patterns and properties of wireless signals [112]. GAI models can identify the presence of jamming signals by continuously observing and analyzing network traffic and can then activate the necessary countermeasures.
To create synthetic samples that closely match real signals, one method is to train GAI models on massive datasets of typical wireless signal patterns. Any appreciable differences between the generated samples and the received signals can be used to identify potential jamming assaults [40]. This improves the responsiveness of security systems by enabling the quick and precise detection of jamming actions. Intelligent jamming mitigation solutions can be developed with the help of GAI. Models can learn to identify and categorize many sorts of jamming attacks by being trained on a wide variety of signals [14]. Using this knowledge, adaptive and robust defenses can be created to dynamically modify their tactics to overcome particular jamming approaches.
Furthermore, jamming attack prediction and prevention can be accomplished through GAI approaches. These algorithms can identify potential weaknesses and forecast the risk of upcoming jamming occurrences by examining past data and patterns [40]. This proactive strategy enables network operators to implement preventive actions to lessen the effects of prospective jamming assaults, such as modifying frequencies or adding security measures. The use of GAI for jamming attack detection and mitigation is not without challenges. These include the need for substantial diverse training datasets, the ability of models to adapt to evolving jamming strategies, and the need for real-time processing for quick response [112]. These issues can be resolved and the defense against jamming assaults in mobile networking can be strengthened with additional studies and developments in GAI.
IV-D Generative Steganography
Data privacy and sensitive information security are of the highest priority for modern wireless networks. To improve data security in mobile networking, generative steganography combines the ideas of GAI and steganography to conceal sensitive information within apparently harmless wireless signals [113]. With the help of steganography, we can hide information such that undesirable recipients cannot detect it. GANs can be used to produce cover signals with secret information. These models study the statistical properties of typical wireless signals and employ this information to include encrypted data within them [114].
In terms of mobile networking security, generative steganography algorithms offer several benefits. They provide a covert way to send sensitive data without prompting suspicion [101]. Generative steganography overcomes usual detection techniques by hiding data within wireless signals that appear to be harmless. Additionally, generative steganography improves mobile networks’ resistance to unauthorized access and data breaches. Since the secret data are integrated into the cover signals, it is safe and impractical to decrypt without proper tools even if attackers intercept the transmission [115].
Furthermore, generative steganography offers a strong defense against assaults meant to compromise the secrecy and authenticity of wireless communications. The process of embedding and extracting concealed information can be extremely efficient and resilient to detection by unauthorized parties using GAI models [116]. However, some issues must be resolved for generative steganography to be used in real-world settings. The challenges include maximizing the capacity of concealed information while minimizing the influence on the cover signals through optimization of embedding algorithms and the technique’s resilience to various sorts of attacks.
Figure 5: Layer-wise representation of a GAI-based IDS for privacy preservation in mobile networking.
Table IV: GAI for Optimizing Wireless Security Enforcement Scenarios.
Apps Ref Model Inputs Outputs Loss function Performance
Sensitive Data Obscurity [117] GAP-GAN Spectrum sharing policies Obfuscated data for adversary’s inference Minimax Balances attacks and utility for spectrum sharing and signal mapping
[118] Federated DL Decentralized traffic statistics Botnet attack detection and anomaly detection results MSE Outperforms standard centrally managed system in attack detection accuracy
[119] Stacking ensemble Darknet traffic data Darknet traffic identification and characterization MSE Robust against four adversarial attacks
Intrusion Detection [120] CVAE Intrusion data Intrusion detection results log-cosh loss term Capable of generating new intrusion data with promising diversity
[121] Improved CVAE Network data features Intrusion detection results Max Better overall accuracy, detection rate, and false positive rate
[106] Autoencoder Heterogeneous network data Intrusion detection results Wasserstein distance Efficiently detects network threats and resolves data imbalance problems
[122] Deep autoencoder Metaverse network traffic data IoT intrusion detection Wasserstein distance Higher accuracy for binary and multiple classification
[105] Transformer Data from 6G-enabled IoT Cyber attack detection results Log discriminator Achieved high overall accuracy in detecting IoT attacks
Jamming Attacks [123] GDNN, GALNN Communication data from cognitive radios Transmission decisions to mitigate jamming attacks Cross entropy Achieves successful transmission by taking optimal actions for defense mechanism
[124] SamBee Frames under different frequencies Jammed ZigBee frames in distinct channels MSE Addresses parallel spoofing attacks and multiple-channels jamming
[125] GAN-enhanced Spectrum sensing information Anti-jamming channel access policy Adversarial & reconstruction Effectively avoids complex jamming attacks
[126] GAN-based Original wireless signal features Probability of misclassification for spoofing attacks Minimax Significant increase in the success probability of wireless signal spoofing
[127] Hidden Markov model Signals processed with sliding window Jamming status of each received energy sample Maximum posterior probability Demonstrates superiority in detecting stealthy jamming without prior knowledge
Generative Steganography [114] GAN based Secret information with noise vector Stegno image containing hidden secret information Minimax Evaluation shows improved image quality and security
[128] GSN Secret and noise data Realistic stego images with secret data Minimax, adversarial loss Superior resistance to steganalysis detection
[34] DCGAN Secret message, uncorrupted region in the image Stego image with hidden secret message Contextual & perceptual Qualitative and quantitative evaluations of the generated stego images.
[129] LSTM Sentence for analysis Prediction of the anticipated texts Cross entropy The optimization algorithm reduces losses in neural networks
[116] GAN, CNN Dealer’s distribution and human face images Revealed message when authorized participants grant consent Noise vector Secure message authentication and multi-factor authentication
IV-E Lessons Learned and Keys Takeaway
Numerous applications of GAI in wireless security have demonstrated promise for bolstering mobile networking defenses against a variety of threats and weaknesses. Differential privacy and safe multi-party computing are two methods that have been successful in obscuring critical information and preserving user privacy. Network security is strengthened against DoS attacks and zero-day vulnerabilities due to the integration of GAI models into network IDS, which enables quick detection of unauthorized access attempts and security breaches. To guarantee network integrity, GAI’s proactive and dynamic intrusion detection algorithms are essential.
In addition, recent developments in GAI have proven crucial in the fight against jamming attempts that interfere with wireless communication. GAI improves malware detection and stops the propagation of threats by analyzing network data and device behavior to quickly identify and isolate infected devices. Also, by encrypting data and disguising it within seemingly innocent material, steganography’s novel application of GAI enables secure data transmission. Concurrently, biometric-based authentication that uses GAI strengthens identity verification processes by recognizing distinctive biometric patterns like fingerprints or facial features. These uses serve as a reminder of the important of GAI for improving data security and facilitating safe and secure user verification.
Mobile networking offers more robust security measures due to the adaptability of GAI in varied security concerns. GAI approaches hold promise to increase the state-of-the-art in mobile network security and communication and data transmission integrity by strengthening defenses against sophisticated attacks. In addition, our review uncovered numerous innovative research methods that further demonstrate the potential of GAI to offer a thorough defence against changing threats and improve the mobile networking security environment, as summarized in Table IV. The future of mobile networking security is bright due to ongoing developments in GAI, which will ensure a safer and more secure digital environment.
In summary, the key takeaways from this section are as follows.
  • GAI strengthens wireless security against anomalies and intrusions in real time.
  • GAI combats jamming attacks and ensures secure data transmission with privacy protection.
  • Biometric-based authentication and malware detection are reinforced using GAI techniques.
  • Ongoing research in this area provides more resilient security measures for mobile networking.
V Application of GAI in Semantic Communication
SemCom revolutionizes communication by addressing limitations faced by conventional systems, even with advanced techniques. It considers the semantics of transmitted data, allowing imperfect decoding while fulfilling the intended purpose and reducing unnecessary data traffic and energy consumption. SemCom presents three primary challenges: exploiting shared knowledge, efficient encoding and decoding, and data aggregation. Overcoming these challenges will be a breakthrough step in offering new alternatives in representation learning and disentanglement learning, accelerating the realization of AI-driven tasks and recommendation systems. Fig. 6 illustrates the general application of GAI in SemCom.
Figure 6: Illustration of GAI in semantic communication. The GAI take place in three blocks, i.e., Knowledge Extractor, DJSCC Encoder and Decoder. The Knowledge Extractor aims to generate both common knowledge, and variant data from the original data. The DJSCC aim to compress and reconstruct the data with a high compression ratio and robust to noise.
V-A Knowledge Abstraction
Knowledge abstraction is crucial for representing abstract knowledge among users and acquiring shared knowledge from distributed data across device domains. KG are the most prominent candidate among contemporary techniques for this purpose. However, manually constructing KGs is labor-intensive and impractical for scalability [130, 131]. Toward automatic KG construction, three FL-enabled designs for GAI deployment in wireless networks have been recently proposed [43]. The success of these design lines is discussed using three key knowledge techniques.
V-A1 Knowledge Summarization
Summarization techniques aim to create concise abstracts that retain essential information for logical reasoning, omitting less crucial details. Therefore, leveraging Language Models (LMs) to extract knowledge from KGs has recently been prevalent. However, automatically learning optimal prompts typically requires abundant training data [132, 133, 134], which may not have useful meaning to new relationships. Despite its success with GPT-3, the approach in [135] is not transferable to other LMs due to its reliance on ad hoc bounded learning and the scalability of GPT-3.
To address the limitations of the aforementioned methods, BertNet [136] proposes an unsupervised approach to automatically generate diverse alternate prompts while requiring the input of minimal relationship definitions. Compared to conventional approaches that rely on extensive human-annotated data or large KGs, BertNet enables extraction of knowledge about new relationships previously inaccessible.
V-A2 Knowledge Representation Optimization
This approach aims to enhance knowledge representation efficiency by transforming complex rule-based structures into more compact formats, such as ontologies or Binary Decision Diagrams (BDDs). A previous study  [137] introduces a spatial-functional embedding for KG completion that represents entity pairs as points and relations as hyper-parallelograms in $\mathbb{R}^{2d}$. This enables pattern representation through the spatial relationship of hyper-parallelograms, providing a clear geometric interpretation of embeddings and captured patterns.
Despite extensive research on knowledge extraction from data, creating representations for relation extraction in a KG remains challenging, especially with distantly labeled datasets due to significant label noise. To tackle this circumvent, a solution called LERP was proposed [138]. LERP is an efficient framework capable of learning prototypes for each relation based on contextual information, capturing the intrinsic semantics of these relations by observing noisy labels.
On the other hand, summarizing KGs encounter two critical problems: limited observations for model training and evolving characteristics of novel entities in temporal KGs. To tackle mose issues, [139] presents MetaTKGR, which employs meta-learning to adjust domain adaptation over time.
V-A3 Knowledge Aggregation
Aggregation techniques combine similar or related knowledge from diverse data sets or learned models to reduce redundancy and strengthen decision-making or the ability to adapt more effectively to new domains [140]. Nevertheless, exploiting automatic knowledge abstraction via entity retrieval encounters three primary obstacles: 1) Context and entity affinity primarily rely on vector dot products, possibly overlooking fine-grained interactions. 2) Large entity sets require a significant memory footprint to store dense representations. 3) Training necessitates subsampling a suitably challenging set of negative data.
Driven by these challenges, a solution called Generative Entity Retrieval (GENRE) is introduced [141] to use a transformer-based generative model in conjunction with a pre-trained NLP model. The unique feature of GENRE lies in its implementation of a constrained decoding strategy, ensuring that every generated entity belongs to a predefined candidate set. As a consequence, the outputs of the GAI model are less likely to deviate from the desired entities.
V-B Data Compression
Data compression is vital in SemCom, particularly for communication with limited resources. The use of lossy compression is believed to be beneficial in improving data compression performance but at the expense of data transmission. With the emergence of SemCom and Goal-Oriented Semantic Communication (GOSC), the shortcomings of lossy compression have been overcome. To achieve this, two main objectives were addressed: improving data representation structure and implementing knowledge-aided techniques. Fig. 7 illustrates the classification of data compression techniques using GAI, potentially strengthening mobile networking robustness.
Figure 7: Taxonomy of data compression techniques through GAI.
V-B1 Structured and Compact Data Representation
In DL and data representation, efficient data compression relies on compact and structured data representation [142, 143]. Recent studies have explored methods in two main categories: optimizing loss functions and designing model architecture.
Loss Function Design. This effort focuses on formulating an efficient loss function to minimize the discrepancies between the original and reconstructed data while capturing noise features from the wireless channel. However, revisiting the classical Shanon theorem [144] indicates that, for a memoryless communication channel, optimal communication involves two distinct processes as the data length approaches infinity:
  1. 1.
    Source Coding: Maximize compression to eliminate redundant information from images.
  2. 2.
    Channel Coding: Employ error-correcting codes to not only facilitate signal reconstruction in the presence of noise but also protect information through redundancy.
While the separation theorem has been demonstrated to be effective in coding algorithms, it encounters two limitations. First, there is no such thing as an infinite number of bits. The overall distortion of signals depends not only on source information but also channel coding errors with finite block length, hindering bit allocation optimization and code design challenges. Second, achieving maximum-likelihood decoding, generally NP-hard, requires impractical computational power, hindering real-world implementation.
Therefore, recent efforts focus on leveraging DL and GAI to JSCC system with acceptable approximation outcomes. Neural Error Correcting and Source Trimming (NECST) is the first effort in this cutting-edge area [145]. This advanced DL framework compresses and corrects input images within a set bit-length constraint. NECST surpasses traditional models by integrating a discrete channel simulation, injecting noise directly into latent representations to enhance resilience in challenging channel conditions. When applied to the Deep JSCC (DJSCC) system [146], NECST activates self-learning, swiftly adapting to new environments without significant computational overheads from training anew. The work in [147] tackles the challenge of JSCC over a discrete-input Additive White Gaussian Noise (AWGN) channel by mapping input images to codewords and decoders to minimize average end-to-end distortion. The specific problem of the end-to-end JSCC is regarded by a quantizer in the middle, where features are extracted from input images and quantized to constellation points through a DNN-based encoder. After transmission, the receiver reconstructs the image from noisy observations using another DNN. The study [148] focuses on mitigating data noise from an impaired channel by capturing noise statistics and applying noise reduction based on channel behavior. Meanwhile, the work in [149] introduces a GAN-based approach for improved data reconstruction in lossy compression. Through a regularization function, this method can address both universal rate distribution and perception functions concurrently.
Error-Bounded Lossy Compression (EBLC) [150] is one critical technique for big scientific data processing that tailors compression according to the precision requirement of scientific datasets and ensures data fidelity within acceptable error bounds. However, to adapt to the highly accurate scientific analysis in diverse conditions, existing EBLC frameworks are mostly limited by configuration issues. To overcome this drawback, the work in [151] introduces a novel framework named OptZConfig to improve accuracy and speed, while the work in [152] develops universal rate distortion-perception representations to enhance data reconstruction.
With the demands of short training time and lower computational power, achieving high compression ratios without compromising data reconstruction, especially in 3D-based AI tasks like point cloud representations, is not a straightforward task. Addressing this, DeepPCAC, a novel approach that directly encodes and decodes point cloud attributes, has been introduced [153]. This method surpasses traditional compressed methods due to adopting geometry instead of using voxelization or point projection.
As previously mentioned, the advent of GOSC signifies the onset of a new era in lossy compression within the field of SemCom, as it does not require reconstructing the entire dataset. Instead, it compresses incoming data as long as AI task performance is achieved. This approach is effective as the AI era enters a prosperous phase, with machine-to-machine and human-to-machine communication becoming more prevalent in IoT Networks [28]. The main feature of GOSC is to extract invariant features among domains [154] or classes [155]. Therefore, as long as the encoders can capture invariant features, they identify crucial AI task features. Based on foundational principles, the authors in [156] integrate invariant learning in a lossy compressor, preserving lossless prediction in AI classification. This method encapsulates essential features for AI tasks, enabling lossy compression while maintaining consistent AI performance.
Architecture Design. The main focus of this effort shifts to the design of an efficient model architecture aimed at enhancing data reconstruction efficiency irrespective of the applied loss function. A typical representative of capturing essential data representations in various media formats is the Attention mechanism, known as Transformer [61]. On this basis, Deep SemCom (DeepSC) architecture for speech recognition applications has been developed to address noise and distortion challenges through joint semantic-channel coding [157]. Similarly, the upgrade versions, U-DeepSC and MUDeepSC [158, 159], accommodate text and image transmission and facilitate semantic information extraction. In that, a layer-wise knowledge transfer technique is exploited to enhance information exchange across tasks, similar to the UNet architecture [160, 161, 162], ensuring detailed information preservation.
Another approach is to design an efficient model architecture for data compression and reconstruction by leveraging the robustness of model diffusion. According to [163], diffusion models comprise three primary components the forward process, the reverse process, and the sampling procedure, which affect the probability distribution of a specified dataset. These models deduce the latent structure of a dataset by capturing the way in which data points propagate through their latent space [164]. Upon the conceptualization of the SemCom process as a sequential Gaussian model [165], the work in [166] introduces an improved lossy compression method capable of seamlessly integrating diffusion models. This method utilizes noise from the SemCom process’s communication channel to improve compression and semantic encoding, enhancing both communication efficiency and data reconstruction noticeably.
V-B2 Knowledge-assisted Compression
Knowledge-Assisted Compression (KAC) is a recent groundbreaking compression approach that inherits from knowledge-based reasoning, as mentioned in Section V-A3. Knowledge-assisted compression approaches tackle the process of compressing or compacting knowledge representations to improve reasoning efficiency and storage requirements. In KAC, systems use explicit knowledge, often represented as rules, facts, or ontologies, to conclude, make inferences, or solve problems. Thus, with the aid of knowledge representation, the compressor and reconstructor can further improve the compression efficiency with the common knowledge among data.
In [167], the authors propose the Residual Compressor (RC) method, akin to knowledge-aided SemCom, where the BPG lossy compression algorithm is used to generate a versatile representation set. This enables subsequent lossless compression and discovers fixed lost information. In addition, a new network architecture with two-stage execution is provided to approximate lost information from lossy compression, which is not only beneficial for reconstructing compressed data but also retains all information. The authors in [168] leverage the generative AI model to transform the original data $S$ into another form of data $\widetilde{S}$. In other words, the generative model $f_{K}(\cdot)$ aggregates the source data with the background knowledge, known by both the transmitters and receivers.
Analogical reasoning is a robust alternative to make reasonable use of KGs in decoding data representation. Fundamental, analogical reasoning aims at increasing machine intelligence by associating data representation (e.g., CV, NLP) with relational, structural, and analogical reasoning. However, research on knowledge-assisted data reconstruction seems to have little focus on knowledge structure. On that basis, a new method for building KGs has been proposed in [169], where Graph Neural Networks (GNN) are considered to develop an efficient data reconstruction technique. Reconstructing data from a single task’s knowledge and data representation may not apply to compressed data operations from multiple semantic knowledge sources. To address this, the work in [170] introduces a competitive embedding framework for logical reasoning over KGs. This approach leverages structure mapping theory and a Transformer-based GAI model to aggregate multimodal knowledge for link prediction, enabling seamless integration into the data reconstruction process. The work in [171] discovers two types of information redundancy in remote sensing image compression: pixel-wise and non-Region-Of-Interest (nROI) data. This allows a more favorable bit allocation strategy by assigning a greater number of bits to the ROI while allocating fewer bits to the nROI. Likewise, the authors in [172] introduce a task-oriented image compression framework that can extract and process ROI and nROI data independently. This framework significantly enhances accuracy and compression efficiency.
To address the oversight of semantic noise and system robustness in data encoding and compression, two efficient methods of Masked Autoencoder (MAE) [173] and Vector Quantized VAE (VQ-VAE) [174] are introduced. Since these methods utilize the GAI model to generate noise and pre-learning adversarial risks before transmission, they not only demonstrate robustness against semantic noise, but also significantly enhance semantic reconstruction via knowledge-based encoder-decoder architectures compared to conventional techniques. The work in [175] presents a quantization-based knowledge extractor, where an automatic mask generation based on semantic knowledge information is developed to regulate the compression ratio of transmitted data. By doing so, the approach enables the SemCom system to dynamically adapt to variable-length coding systems. To further improve compression efficiency within the SemCom system introduced in [157], the authors of [176] employ LSTM techniques to formulate a memory-aided SemCom system. With historical information about the data and compression efficiency, this approach yields higher compression efficiency without compromising communication efficiency, meaning there is no need to transmit additional data via the communication channel.
The research in [177, 178] introduces a new SemCom framework aimed at reliable communication with minimal data transfer across diverse services and users. Central to this framework are GFlowNets, inspired by GAI networks, which enable causal reasoning for SemCom. These networks serve as causal describers to highlight transitions in task distributions with changing contexts when using GFlowNets to identify task event transitions and form a graph for GNN to improve data reconstruction. The study in [179] introduces a causal semantic communication method for abstracting knowledge to enhance communication efficiency among distributed clients. In that, a generative AI model is deployed with the ability to produce causally invariant representations, facilitating the SemCom system to prioritize transmission of the most meaningful features to the goal-oriented task at the recipient’s end. The research in [180] proposes a knowledge-enhanced SemCom framework in which the receiver can more actively utilize the data in the knowledge base for semantic reasoning and decoding, based on only affecting the parameters rather than the structure of the neural networks at the transmitter. This is achieved using a transformer-based knowledge extractor to find relevant factual triples for the received noisy signal.
V-C Generative AI and Distributed Learning
GAI has demonstrated remarkable robustness in enhancing performance in distributed learning. To be specific, GAI techniques produce generative data representations that extract generalized knowledge or meaningful features, supporting training on distributed local devices without requiring access to unseen data domains.
Figure 8: Architecture of integration of GAI into FedGen.
In FedGen [181], the server based on local model performance cooperatively trains a GAI model as a lightweight generator to amass user information and that is capable of generating pseudo data representation without accessing their true data (see Fig. 8). Thus, each local user can have the additional knowledge of other distributed clients, without sharing data among distributed devices. This approach is robust as this innovative approach can significantly enhance the performance and communication efficiency of FL, all while ensuring that user data remain secure and private. In [44], an AaaS framework based on AIGC models is designed to address the challenges in training and deployment overheads for edge networks.
Similar to FedGen, which leverages GAI to train a generative AI model, DENSE [182] designs a GAN-based data generation network that enables the generation of on-server data. In contrast to FedGen, where GAI is employed for on-server data generation, DENSE employs a multi-teacher knowledge distillation technique (see Fig. 9). This technique ensure that the generated data accurately captures the entire system’s data distribution and characteristics. The local models collaboratively instruct a student, denoted as the global model, inheriting knowledge from all teachers. Furthermore, this approach can be enhanced by incorporating Label-Driven Knowledge Distillation (LKD) [183]. This methodology enables the student to selectively learn the most meaningful knowledge from each teacher, significantly improving distillation efficiency compared to other conventional knowledge distillation approaches. Consequently, it is feasible to apply the ensemble learning on the server, boosting the FL efficiency while retaining the communication efficiency.
Different from other GAN approaches, the work in [184] adopts a distinct approach by treating model parameters as data. Specifically, the authors introduce HCFL, a technique capable of compressing and reconstructing model parameters. It enables the FL system to circumvent substantial communication overheads. In particular, applying HCFL to FL systems can maintain stable accuracy when the distortion rate of the aggregated model decreases proportionally with network size.
Figure 9: Architecture of integration of GAI into DENSE.
Aside from the aforementioned methods, invariant learning is a unique approach to enhancing learning on local devices in distributed learning. In invariant learning, users collect their local invariant representations or cross-domain invariant representations by sharing the most consistent format of data (invariant features) across devices. In this way, it minimizes the need for extensive communication resources, a representative form of knowledge-aided SemCom in distributed learning. For instance, Federated Cross-Correlation and Continual Learning (FCCL) [185] and FCCL+ [186] are innovative methods inspired by Barlow Twins principles, which help identify invariants across domains in the FL system (See Fig. 10). As an alternative method to discover invariant features, the authors in [187] focus on class-wise invariance by introducing cluster prototypes of contrastive learning. Unlike [185, 186, 187], FedPAC employs a structured causal model to acquire goal-oriented invariant representations, demonstrating robustness in learning personalized representations for each client without extensive knowledge sharing among devices [188].
Figure 10: Architecture of integration of GAI into FCCL.
Table V: GAI for Optimizing Semantic Communication Scenarios.
Apps Ref Model Inputs Outputs Loss function Performance
Knowledge Abstraction [141] Transformer-based GAN Textual input source Summarized knowledge Cross Entropy Competitive results on document retrieval with less memory
[136] BERT GAN Language Data Summarized Knowledge Log-likelihood Only necessitates a basic definition of relations as input, making it appropriate for extracting knowledge about novel, intricate relations that are immediately established and were previously unavailable
[189] LERP Knowledge Graph with missing edges Contextual information as vectors Gradient-based Outperforms other rule learning and estimates expressive logical rules
[190] ADT-GAN Multivariate time series data Anomaly scores for each data point Adversarial Loss Effectiveness in anomaly detection within multivariate time series data
 [22] LLMs Data and information from edge networks Collaborative planning and decision-making by multi-agent LLMs MSE Demonstrates the potential benefits of on-device LLMs to solve network intents.
 [191] DGMs Distorted reconstructions of images Improved perceptual quality of transmitted images MSE Improves perceptual quality by performing denoising using style-based GAN
Data Compression  [192] VQ-VAE Spectral envelope parameters from speech signals Improved quantized spectral envelope parameters for speech vocoder Vector quantization Reduces log spectral distortion and increases perceptual evaluation of speech quality
 [193] AE-LSTM Heterogeneous data from IoT devices Identification of out-of-norm activities for cyber threat hunting in IIoT MSE Ensemble model outperforms conventional ML with better compression
 [94] ELAN Captured images from edge devices Semantic optimized resource allocation extracted from images MSE Reduces transmission costs by extracting important semantic information
 [148] NECST Heterogeneous data from IoT devices Improved quality reconstructed images MSE Significant gain in data reconstruction under different settings of channel distortion
 [157] DeepSC Captured images from edge devices Realistic image reconstructed based on semantic segmentation BLEU Immense resource gain in bandwidth reduction during semantic segment transmission
 [167] RC Captured images from edge devices Lossless reconstructed images. Cross-entropy Lossless image reconstructed from lossy compressed data
 [179] Causal semantic communication Sequential signal Reconstructed sequential signal KL Divergence Improves signal reconstruction using causal theorem to predict the potential signal to support the reconstruction
 [177, 178] NeSy Graph representation Knowledge based on graph KL Divergence Using neural-symbolism to improve the AI task of reconstructing knowledge from the graph
Semantic Resource Allocation  [194] DRL Real-time data from vehicular networks Improved service quality, reliable guidance to receiving vehicles MSE Enhances service quality, reliability, and efficiency in GAI-enabled vehicular networks
 [195] ISGC User inputs for semantics and graphics rendering Optimized resource allocation and high-quality content generation Variational autoencoder Optimal resource allocation strategy for the Metaverse
 [196] Diffusion Wireless network collaboration tasks Optimized execution of AIGC tasks MSE Optimized edge computation resource utilization
 [197] RSA Multi-modal data from human/machine agents Locally rendered scenes and interactions in the Metaverse MSE Utilizes distributed learning and multi-agent RL for efficient communication
Distributed Learning  [22] LLMs Edge networks and multi-agent system parameters Collaboratively planned and solved tasks in wireless networks Custom ensemble loss Enables self-governed networks with on-device LLMs for intelligent decision-making
 [43] FL-GAN AIGC tasks, settings of wireless networks, and privacy Diverse, personalized, and high-quality content generation Adversarial loss Improved learning efficiency, privacy protection, and content quality
[181] FL-GAN Heterogeneous user data Ensembled knowledge for guiding local models KL Divergence Improved generalization performance in FL with minimum communication resource
 [198] Diffusion Network scale, environment parameters, and user demands Customized network solutions and expert-free optimization Custom ensemble loss Best adaptation to objectives and scenarios and creation of novel designs
V-D Lessons Learned and Keys Takeaway
GAI plays a crucial role in the development of SemCom, especially in generating data suitable for various subject-specific requirements. Consequently, applying GAI to various subdomains brings several benefits to SemComs, including creating more compact and structured data representations, significantly improving data compression efficiency in SemCom, and enhancing communication efficiency.
Furthermore, the robust development of GAI, particularly in summarization and knowledge extraction, enables the system to extract common knowledge from data. In other words, identifying elements that are domain-invariant or label-invariant (consistent regardless of changes in data) allows the system to leverage such invariant information multiple times without passing it through the physical communication channel. This leads to a substantial reduction in communication overhead without diminishing the amount of transmitted information.
In summary, the key takeaways from this section are as follows.
  • GAI markedly enhances compression performance through the creation of concise and organized data representations.
  • GAI empowers the SemCom capability to abstract knowledge effectively, facilitating high-performance data compression without sacrificing information.
  • GAI facilitates the generation of pseudo data on distributed devices, allowing these devices to adapt to the domains of other devices in distributed learning scenarios.
  • GAI can assist the encoder in generating invariant features from data, comprising the most meaningful features that contribute significantly to AI performance. This improvement leads to enhanced learning while concurrently achieving a substantial reduction in communication costs.
  • While current knowledge abstraction heavily relies on NLP models, there exists significant research potential in SemCom. There is a pressing need for the introduction of a generalized, scalable approach to abstract knowledge across types, eliminating the necessity of intermediate processes like NLP.
VI Challenges and Future Directions
GAI has demonstrated its effective abilities in the development of mobile networks, such as enhancing the user experience, personalizing services, optimizing network performance, and enabling new business models. In return, it also poses several new issues that need to be addressed. In what follows, we will explore and propose numerous viable solutions to technical challenges of GAI (Section VI-A), including lack of interpretability (Section VI-B); ethical and legal concerns (Section VI-C); and security, privacy, and trust (Section VI-D). Fig. 11 illustrates the categories of open research challenges and directions prevalent in mobile networking through GAI. Finally, we summarize all critical issues and the respective solutions in Table VI.
Figure 11: Open research directions in mobile networking through GAI.
Table VI: Summary of Challenges, Critical Issues, and the Respective Solutions of GAI in Mobile Networking
Challenges Issues Potential Impact on Mobile Networking Applications Future Research Directions
Handling technical complexity Improved performance and cost trade-offs in GAI-based mobile networks Develop resource-efficient GAI models
Technical Challenges Tackling legacy system Integration challenges and security risks Explore SDN and network slicing integration
Avoiding technical debt Improved quality and reliability of GAI models Implement continuous learning and improvement
Potential misuse and AI hallucinations Generation of harmful or biased content Develop content authentication and moderation methods
Data handling violations Protection of personal data and privacy rights Implement data anonymization and encryption
Human-AI interaction Ethical and social implications in mobile network workforce Explore human-in-the-loop approaches
Integration challenges Compatibility issues with existing infrastructure Develop new network paradigms and standards
Scalability Resource limitations for large-scale GAI deployment Explore cloud and edge computing solutions
Complexity Difficulty in interpreting complex GAI models Focus on transparent and interpretable algorithms
Lack of Interpretability Diversity Challenges in comparing and integrating diverse data sources Utilize visualization for diverse data
Uncertainty Difficulty in dealing with probabilistic and stochastic models Employ quantification for uncertainty estimation
Unintended consequences Unforeseen outcomes from GAI decisions Develop interpretability metrics and evaluation
User trust Lack of user trust due to opaque decisions Enhance interpretability for user confidence
Authenticity Risk of fake or misleading content generation Develop digital watermarking and verification methods
Ethical and Legal Concerns Ownership Determining ownership and intellectual property rights Implement digital attribution and licensing systems
Fairness Bias and unfairness in content generation Incorporate fairness mechanisms
Responsibility Accountability and responsibility for GAI-generated outcomes Integrate responsibility by design principles
Societal impact Broader ethical implications on society Address broader ethical considerations in GAI systems
Unauthorized access Breach of system security Implement strong authentication and access controls
Security, Privacy, and Trust Data leakage Unauthorized sharing of sensitive information Enhance data encryption and secure communication
Trustworthiness assessment Difficulty in evaluating GAI reliability Develop trust evaluation frameworks and metrics
Adversarial attacks Targeted attacks on GAI models Enhance security against adversarial attacks
VI-A Technical Challenges
To enable new services and use cases for future mobile networks (i.e., SemCom, mobile edge computing, user behavior analysis, wireless signal processing, channel rendering, 3D mapping, and integrated sensing), GAI faces some technical challenges that need to be addressed. The details of each challenge and the corresponding solution are discussed in the following.
Handling technical complexity: GAI models often require powerful computing ability and data to train and fine-tune, necessitating mobile network components to sustain the trade-off between performance and cost. For instance, ChatGPT-3, a text generation tool, requires 175 billion parameters and hundreds of billions of words from the internet. In this context, exploiting cloud or edge computing services is crucial to reducing computation and storage processing tasks from mobile devices or networks. Also, integrating model compression approaches into GAI systems like pruning, quantization, or distillation are also needed to reduce model size and complexity without affecting performance.
Tackling the legacy system: The incompatibility of GAI models with existing network infrastructure and protocols can limit their integration and scalability. In particular, current mobile networks, developed based on centralized servers and fixed bandwidth resources, may not work well with the GAI model and can even generate risks involving network security, privacy, and reliability. To cope with these issues, there is a need for new network paradigms such as SDN and network slicing to provide more flexibility, efficiency, and customization for GAI applications. In addition to this, it is also necessary to develop new standards and regulations to ensure the interoperability of GAI models on mobile networks.
Avoiding technical debt: The risk of falling behind the latest developments and best practices in GAI can affect the quality and reliability of services. Mobile network operators must keep up with the latest research and innovation in the field and update the deployed AI models to avoid errors, bugs, or vulnerabilities that may arise over time. Following that, two possible solutions should be considered: (1) employing a continuous learning and improvement approach to monitor the performance and feedback of GAI models and to make adjustments accordingly and (2) collaborating with other stakeholders such as researchers, developers, or customers to share general AI knowledge and best practices.
Potential misuse and AI hallucinations: GAI models without control limitations can compromise the quality and reliability of content, potentially producing fake news, spam, phishing emails, and malware. Attackers can exploit these to manipulate users or compromise their security. Hence, feedback and evaluation mechanisms, such as digital watermarking or blockchain, are needed to verify the authenticity of GAI content, along with moderation or filtering techniques to detect harmful content. To further prevent security threats related to the generation of adversarial examples, GAI-based systems must have strong and flexible security measures, such as encryption, authentication, anomaly detection, and prevention. In particular, GAI models may require access to personal data without proper consent or protection, resulting in sensitive information leaks. To ensure data privacy, GAI-based mobile systems need to have mechanisms to encrypt, anonymize, or delete data after use.
Human-AI interaction: The integration of GAI into mobile networks to perform/replace some tasks can impact the skills and roles of the workforce, as well as the ethical and social aspects of employment, education, training, and motivation. This raises questions about GAI models’ accountability, transparency, fairness, and trustworthiness. To address this, human-in-the-loop or human-on-the-loop approaches are needed to create a tight connection between human workers and GAI in terms of design, operation, and evaluation. At the same time, building ethical principles and guidelines is critical to ensure the responsible and beneficial use of GAI in mobile networks as well as the shared values and norms of users and providers.
VI-B Lack of Interpretability
Lack of interpretability, another challenge of GAI for mobile networks, refers to lack of explaining and understanding the reasons behind GAI-based system outputs, particularly important in critical or sensitive tasks like network security, performance optimization, or service personalization. Interpretability acts as a bridge between mobile users and providers, fostering trust and confidence. For example, errors, failures, or unexpected outputs produced by GAI-based systems (i.e., generating harmful or biased content), interpretability becomes a powerful tool to help users and providers debug and improve the quality of GAI models. When there are growing concerns regarding ethical or legal issues (e.g., data violations, copyright infringement, or misinformation), interpretability will be the basis to assist users and providers in solving accountability and responsibility problems. It is not straightforward to obtain these benefits due to the following reasons.
Complexity: GAI-based systems are often developed based on complex and non-linear algorithms. For example, DNN contains a lot of hidden layers, parameters, and interactions that are difficult to interpret or visualize. Hence, GAI-based systems should adopt simpler or more transparent algorithms, such as linear models, decision trees, or rule-based methods.
Diversity: As the inputs of GAI models are diverse and heterogeneous (i.e., text, audio, video, or code), it is hard to perform comparison and integration due to their different formats, structures, and even semantics. A possible solution to overcome such challenges is to exploit visualization features to enhance the diversity of GAI-based systems due to graphical and interactive tools (e.g., charts, diagrams, or dashboards), which allow analysis and visualization of the data or outputs of GAI models more intuitively or understandably.
Uncertainty: Most GAI models rely on probabilistic and stochastic models to produce estimated outputs, such as GANs. Therefore, the randomness, variability, and noise during the training process complicate the measurement, processing, and control. The use of quantification features, in this case, is a reasonable choice to mitigate the uncertainty of GAI-based systems, e.g., numerical or statistical approaches such as metrics, scores, or confidence intervals, as well as in estimating the accuracy and reliability of the outputs.
VI-C Ethical and Legal Concerns
Ethical and legal concerns arise when applying GAI to mobile networks as they relate to the ethical and legal implications of creating, using, and sharing new content. These may affect the rights, interests, values, and responsibilities of network stakeholders, such as developers, users, operators, and regulators; the law; and society in general.
Authenticity: In mobile networking, GAI-generated content introduces concerns regarding the authenticity and reliability of the information. This requires measures like digital watermarking to establish source identity and origin identification as well as digital verification techniques (i.e., cryptography or statistics) to ensure content integrity. More importantly, mobile network users need to be further educated with digital literacy-oriented programs to have a basis for evaluating and verifying GAI-generated content and its sources.
Ownership: The emergence of creative content from GAI systems raises concerns about ownership, authorship, and intellectual property rights in mobile networks. Digital rights management solutions become essential to protect and manage these rights, ensuring fair use and preventing unauthorized distribution. GAI-based systems should incorporate digital attribution mechanisms to acknowledge the contributions of original data sources from wireless networks and mobile devices. Furthermore, the development of digital licensing systems is crucial for specifying and enforcing terms and conditions governing the use, modification, and distribution of GAI-generated content.
Fairness: GAI systems raise concerns about bias in the contents generated from data sets, algorithms, or even with creators. Ensuring fairness characteristics of diversity and inclusion requires long-term investments during the deployment of GAI systems for mobile networks. In that, implementing fairness mechanisms, including auditing, debiasing, and calibration, becomes crucial to detect and rectify biases or errors. In addition to this, establishing and enforcing fairness policies are vital steps in defining rules and obligations for GAI systems in active mobile networking environments.
Responsibility: Impactful information generated by GAI-based systems raises concerns about responsibility and accountability. Determining who is liable for the outcomes of GAI systems and how to ensure adherence to ethical and legal standards becomes paramount, particularly for mobile networking applications. Addressing these concerns is not limited to responsibility mechanisms (e.g., feedback, reporting, and review processes) to monitor and assess the impacts of GAI systems but also involves cooperation between developers and users in establishing policies to uphold ethical norms and regulations in the mobile networking context.
VI-D Security, Privacy, and Trust
Although GAI gains significant enhancements in the performance and efficiency of mobile networks, there are various challenges and risks associated with integrating GAI into mobile networks, especially in terms of security, privacy, and trust. In what follows, we will discuss the challenges and the respective solutions for these issues.
VI-D1 Security
Security refers to the ability to protect GAI-based systems and the data they use or produce from unauthorized access, modification, and damage. In particular, GAI can be vulnerable to cyberattacks, such as hacking, spoofing, DoS, or adversarial learning. These attacks can compromise the integrity, availability, and confidentiality of the network and its data. In particular, when GAI is poisoned, it becomes a powerful tool for attackers to launch attacks on other systems/users or to manipulate/extract sensitive information.
To ensure security in mobile networks with GAI, some measures could be taken. First, there is a need to establish security policies and rules as the basis for operating, updating, and maintaining the safety of GAI-based systems. Second, it is required to implement security mechanisms to protect GAI systems from external or internal attacks, such as encryption, authentication, authorization, firewalling, monitoring, and auditing. Third, there is a need for security testing to evaluate and verify the security performance and resilience of GAI systems under different scenarios and conditions. Finally, security awareness and education are needed to raise the knowledge and skills of GAI users and operators on how to prevent and respond to security incidents.
VI-D2 Privacy
Privacy refers to the ability to respect and protect the personal or sensitive information of users. Specifically, the integration of GAI into mobile networks could pose a risk to user data privacy. In that, GAI could collect, process, store, or share personal or confidential data without proper consent, transparency, or accountability. Moreover, GAI could leak sensitive information from the data, such as identity, location, behavior, preferences, or health status. This could violate the privacy rights and expectations of the network stakeholders.
To cope with the privacy issues in GAI-empowered mobile networks, some potential measures should be considered. First, it is required to use privacy mechanisms to safeguard personal or confidential data, such as anonymization, pseudonymization, encryption, access control, and data minimization. Second, there is a need to establish policies and rules to maintain user privacy in GAI systems. Third, privacy tests are necessary to reduce the potential privacy risks and harms that could be caused by GAI systems. Finally, privacy notices and choices is needed to inform the network stakeholders about their privacy rights and options.
VI-D3 Trust
Trust refers to the ability to ensure the reliability, accuracy, and quality of GAI-based systems and the outputs they generate. Normally, GAI could undermine the perception of reliability, competence, honesty, and benevolence of a system or a provider by being unpredictable, inconsistent, and opaque. For example, GAI could make errors or mistakes that affect the quality or availability of the network services. In addition, GAI could hide or falsify its actions that affect the security or privacy of the network.
To tackle such problems, some potential solutions can be exploited for GAI-based systems. First, there is a need to deploy trust mechanisms for GAI systems and their interactions, such as reputation, feedback, certification, verification, and explanation. Second, policies and rules are needed to maintain the trustworthiness of GAI systems to users. Third, a greater focus on measuring and monitoring trust assessments across diverse criteria and metrics will help increase the reliability and performance of the GAI system. Finally, trust education and communication programs are needed to enhance the credibility of GAI systems and the benefits of stakeholders.
VII Conclusion
In this article, we comprehensively reviewed the fundamentals and potential applications of GAI for the future Internet and mobile networking. We first provided the evolution and principles of GAI and then discussed its applications for important research issues in mobile communications and networking, including network management, wireless security, and SemCom. In each application, we provided lessons learned and key takeaways from the reviewed literature to help researchers develop research into applied GAI for mobile networking. As research on GAI for mobile communications and networking is in the early stages, we need to consider both opportunities and challenges toward the realization of GAI in this research area. As such, we finally discussed key challenges and potential research directions, such as technical challenges, privacy, security, trust, ethical and legal concerns, and open-source frameworks and tools.